# Rule-based Generation of de Bruijn Sequences: Memory and Learning

### v1

@@ Abstract
We investigate binary sequences generated by non-Markovian rules with memory length $\mu$, similar to those adopted in Elementary Cellular Automata. This generation procedure is equivalente to a shift register and certain rules produce sequences with maximal periods, known as de Bruijn sequences. We introduce a novel methodology for generating de Bruijn sequences that combines: (i) a set of derived properties that significantly reduce the space of feasible generating rules, and (ii) a neural network-based classifier that identifies which rules produce de Bruijn sequences. Experiments for large values of $\mu$ demonstrate the approach’s effectiveness and computational efficiency.
@@ 1 Introduction
Cellular Automata (CA) are a class of dynamical systems that evolve in discrete time and space (Wolfram1983, ; Wolfram1984, ). A CA consists of cells (or agents), each of which can adopt a state from a discrete set. The most commonly used alphabet is binary, typically denoted by $\{0,1\}$. The CA evolves by applying generating rules at each time step. These rules can be static or dynamic (i.e., they may change depending on the current state of the entire CA), and either local or global, depending on whether they involve only neighboring cells or the entire population.
@@
Relatively less attention has been given to zero-dimensional cellular automata (0dCA), which consist of a single cell whose state is updated based solely on its own past states. Despite their apparent simplicity, 0dCA do not involve interactions between multiple cells, their time evolution yields interesting sequences of symbolic states. The rules governing such systems depend on the memory $\mu$—the number of previous states influencing the next state—and the alphabet $\{\alpha\}$, the number of possible states a cell can take. Generally, the complexity of the system increases with both $\mu$ and $\alpha$. For simplicity, we focus on binary systems, i.e., $\alpha=2$, using the symbols $\{0,1\}$.
@@
Symbolic sequences naturally arise in discrete dynamical systems, e.g. CA (Jin, ). From an initial condition, the next state is computed according to a generation rule (or function) that depends on previous states. For instance, a classical non-Markovian binary time series can be defined by initial values $a_{1}$ and $a_{2}$, and a recursive rule:
@@
With initial states $a_{1}=0$ and $a_{2}=1$, the sequence continues as $a_{3}=1$, $a_{4}=0$, $a_{5}=1$, $a_{6}=1$, etc., producing the binary sequence:
@@
Other examples with $\mu=2$ can be constructed using logical operators. For instance, applying the AND function with the rules:
1 11,1 00,0 10,0 00formulae-sequence111formulae-sequence100formulae-sequence0100001\,1$\rightarrow$ 1$\quad$ 1\,0$\rightarrow$ 0$\quad$ 0\,1$\rightarrow$ 0$\quad$ 0\,0$\rightarrow$ 01 1 → 1 , 1 0 → 0 , 0 1 → 0 , 0 0 → 0
to the initial pair 0 1010\,10 1 produces the sequence:
@@
Similarly, applying the OR function:
1 11,1 01,0 11,0 00formulae-sequence111formulae-sequence101formulae-sequence0110001\,1$\rightarrow$ 1$\quad$ 1\,0$\rightarrow$ 1$\quad$ 0\,1$\rightarrow$ 1$\quad$ 0\,0$\rightarrow$ 01 1 → 1 , 1 0 → 1 , 0 1 → 1 , 0 0 → 0
to the same initial configuration 0 1010\,10 1 results in:
@@
In contrast, applying the XOR function:
1 10,1 01,0 11,0 00formulae-sequence110formulae-sequence101formulae-sequence0110001\,1$\rightarrow$ 0$\quad$ 1\,0$\rightarrow$ 1$\quad$ 0\,1$\rightarrow$ 1$\quad$ 0\,0$\rightarrow$ 01 1 → 0 , 1 0 → 1 , 0 1 → 1 , 0 0 → 0
to the initial state $a_{1}=0,a_{2}=1$ yields:
[…]
with period $T_{4}=3$. However, using the same rule with the initial configuration: $a_{1}=0,a_{2}=0$ results in a constant sequence $a_{k}=0$ for all $k>2$. These examples illustrate that the resulting behavior depends not only on the rule but also on the initial conditions.
For any finite memory $\mu<\infty$, all sequences generated in this framework are eventually periodic. That is, for any rule and initial condition, the sequence will eventually settle into a repeating pattern (also referred to as a motif or loop). The maximum possible period for a sequence with memory $\mu$ is:
@@
As $\mu$ increases, the sequences may resemble aperiodic ones, making them useful for computing pseudorandom numbers to be applied in Monte Carlo simulations, and cryptographic systems (Etzion, ). Truly aperiodic behavior can only arise when $\mu\rightarrow\infty$, or when the alphabet is infinite (e.g., using real-valued states).
@@
where $a_{i}\in\{0,1\}$ for all $i$ (see Fig. 1). Starting from an initial configuration, the rule is applied recursively to generate the digits of the sequence as follows:
@@
For example, when $\mu=3$, there are $2562^{2^{3}}=$ 256 possible rules—corresponding to Wolfram’s elementary 1dCA rules. Figure 1A shows the XOR rule for $\mu=3$, represented by the binary rule string 10010110, which corresponds to rule number 150 in Wolfram’s classification. Applying this rule to the initial configuration 0 1 00100\,1\,00 1 0 yields:
@@
In general, for a given $\mu$, the total number of possible generating rules is $2^{2^{\mu}}$, and the maximum number of binary sequences generated from these rules is:
$C(\mu)=2^{2^{\mu}+\mu}$
This number grows super-exponentially with $\mu$, making exhaustive analysis computationally infeasible for large $\mu$. For instance, when $\mu=10$, there are approximately $10^{308}$ rules, each of which must be applied to 1024 different initial configurations.
This raises the fundamental question: given a memory value $\mu$ and an initial configuration, is it possible to predict the asymptotic pattern and period $T$ of the resulting sequence? While this is tractable for small $\mu$ via exhaustive enumeration, alternative methods are necessary for large $\mu$ due to the sheer size of the rule space.
Figure 1: Two generating rules with memory $\mu=3$. A) Rule 150, whose binary representation is $10010110$. This corresponds to the following assignments for each 3-tuple (ordered from $111$ to $000$): $111\rightarrow 1$, $110\rightarrow 0$, $101\rightarrow 0$, $100\rightarrow 1$, $011\rightarrow 0$, $010\rightarrow 1$, $001\rightarrow 1$, $000\rightarrow 0$. The application of this rule to four different initial triplets is shown on the left. As observed, the resulting sequences are periodic with periods 1, 4, 1, and 2 (from top to bottom). The dynamics of this rule can be visualized as a directed graph, shown at the bottom. Each node represents a possible 3-tuple; a directed edge from node $x$ to node $y$ exists if applying the rule to $x$ yields the last digit of $y$. This graph is not connected and does not contain a Hamiltonian path. B) Rule 45, with binary representation $00101101$, corresponds to the truth table shown on the left of panel B. In contrast to Rule 150, applying Rule 45 to any of the $82^{3}=$ 8 initial triplets yields the same cyclic sequence of maximum period $T_{\text{max}}=8$, namely, the de Bruijn sequence $s_{6}=00010111$. This rule is thus a de Bruijn rule. The associated graph, built as described above, is a de Bruijn graph because it contains a Hamiltonian path through all $2^{3}$ nodes.
It is worth noting that the 0-dimensional Cellular Automata (0dCA) with memory, as defined above, are equivalent to sequences generated by a Linear Feedback Shift Register (LFSR). This type of sequence has been extensively studied, and a vast body of literature exists on the subject (Golomb, ). In particular, significant attention has been devoted to the characterization and computation of shift register sequences with maximum period—commonly known as de Bruijn sequences (deBruijn, ; Fredricksen, ; Ralston, ). As we will show in the next section, there exists a one-to-one correspondence between de Bruijn sequences and their generating rules, which we will refer to as de Bruijn rules.
@@ 3 Rules that generate maximum period sequences: de Bruijn rules
For a given memory length $\mu$, the maximum period that a generated binary sequence can attain is $T_{\text{max}}=2^{\mu}$. Naturally, the minimum period is 1, corresponding to sequences that converge to fixed points (e.g., either 0 or 1). For small values of $\mu$, it is feasible to generate all possible sequences and compute their corresponding periods. Figure 2 and Table 1 show, respectively, the bar chart of the observed periods and the associated frequencies for the case $\mu=4$.
Table 1: Number of rules that generates sequences with each of the possible periods that can appear, from 1 to 16, for $\mu=4$. The bar chart exhibits three peaks at $T=1$, $T=3$ and $T=5$. For $T>5$ the distribution decreases monotonously (see Fig. 2)
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
16784 7220 9547 8060 8665 5668 3494 2670 1592 977 421 192 90 77 63 16
Figure 2: Bar chart of periods for a memory $\mu=4$. The frequencies correspond to the values presented in Table 1. Similar bar charts appear for other $\mu$-values. Note that the maximun period has the lower occurrence; as a matter of fact, there are $2^{2^{3}-4}$ de Bruijn sequences.
@@
Sequences with maximum period, known as de Bruijn sequences, are of particular interest. These are cyclic sequences of length $2^{\mu}$ in which every possible binary substring of length $\mu$ appears exactly once. For small values of $\mu$, it is straightforward to generate all possible de Bruijn sequences. There exists a unique de Bruijn sequence for $\mu=2$, namely 0011. For $\mu=3$, there are two such sequences: 00010111 and 00011101, that are generated by rules 45 and 75, respectively (rule 45 is graphically described in Figure 1B).
@@
Since these sequences are cyclic, their representations are not unique; by convention, we select the lexicographically least sequence (i.e., the smallest in binary order and decimal number) as the representative of the equivalence class. Specially significative is the lexicographically least sequence for each memory $\mu$, that is referred to as the granddaddy in Knuth .
@@
The number of binary de Bruijn sequences for a given $\mu$ is given by $C(\mu)=2^{2^{\mu-1}-\mu}$ (deBruijn, ; A016031, ). This number grows also superexponentially with $\mu$ and reaches extremely large values even for moderate $\mu$. For example, for $\mu=10$, $C(10)\approx 10^{151}$. Generating all de Bruijn sequences is a challenging combinatorial problem that remains unsolved in full generality, though various efficient construction algorithms have been proposed (see, for instance, deBruijnproject ).
@@
Table 2: Correspondence betweeen de Bruijn rules and de Bruijn sequences for $\mu=1,2,3$ and $4$. It is also shown the Evil Odd Number that divides each de Bruijn rule, leaving the remainder $\phi(\mu)$.
@@
4 43 255 10965 0010101011010101 0000101001101111
4 57 255 14535 0011100011000111 0000110111100101
4 65 255 16575 0100000010111111 0000111101001011
4 71 255 18105 0100011010111001 0000100111101011
4 75 255 19125 0100101010110101 0000101111010011
4 83 255 21165 0101001010101101 0000101100111101
4 85 255 21675 0101010010101011 0000111101011001
4 89 255 22695 0101100010100111 0000110010111101
4 99 255 25245 0110001010011101 0000101001111011
4 113 255 28815 0111000010001111 0000111101100101
For $\mu=4$, there are $C(4)=2^{2^{3}-4}=16$ de Bruijn rules and their associated de Bruijn sequences. As it can be seen, the granddaddy sequence is: $0000100110101111$. The corresponding granddaddy rule is $3825$, with binary representation $0000111011110001$. For $\mu=5$, the number of de Bruijn sequences is $C(5)=2^{2^{4}-5}=2048$. The granddaddy rule in this case is $00001100111111101111001100000001$ (whose decimal representation is 218034945), which generates the granddaddy sequence: $00000100011001011011101010011111$.

### v2

@@ Abstract
We investigate binary sequences generated by non-Markovian rules with memory length $\mu$, similar to those adopted in elementary cellular automata. This generation procedure is equivalent to a shift register, and certain rules produce sequences with maximal periods, known as de Bruijn sequences. We introduce a novel methodology for generating de Bruijn sequences that combines (i) a set of derived properties that significantly reduce the space of feasible generating rules and (ii) a neural-network-based classifier that identifies which rules produce de Bruijn sequences. The experiments for some values of $\mu$ demonstrate the approach’s effectiveness and computational efficiency.
@@ 1 Introduction
Cellular automata (CA) are a class of dynamical systems that evolve in discrete time and space (Wolfram1983, ; Wolfram1984, ). A CA consists of cells (or agents), each of which can adopt a state from a discrete set. The most commonly used alphabet is binary, typically denoted by $\{0,1\}$. The CA evolves by applying generating rules at each time step. These rules can be static or dynamic (i.e., they may change depending on the current state of the entire CA) and either local or global, depending on whether they involve only neighboring cells or the entire population.
@@
Relatively less attention has been given to zero-dimensional cellular automata (0dCA), which consist of a single cell whose state is updated based solely on its own past states. Despite their apparent simplicity, 0dCA do not involve interactions between multiple cells, their time evolution yields interesting sequences of symbolic states. The rules governing such systems depend on the memory $\mu$—the number of previous states influencing the next state—and the alphabet $\{\alpha\}$, the number of possible states a cell can take. Generally, the complexity of the system increases with both $\mu$ and $\alpha$. For simplicity, we focus on binary systems, i.e., $\alpha=2$, using the symbols $\{0,1\}$, and, due to computational limitations, we restrict our study to a subset of cases with relatively low values of $\mu$ (less than 8).
@@
Symbolic sequences naturally arise in discrete dynamical systems, e.g., CA (Jin, ). From an initial condition, the next state is computed according to a generation rule (or function) that depends on previous states. For instance, a classical non-Markovian binary time series can be defined by initial values $a_{1}$ and $a_{2}$ and a recursive rule:
@@
With initial states $a_{1}=0$ and $a_{2}=1$, the sequence continues as $a_{3}=1$, $a_{4}=0$, $a_{5}=1$, $a_{6}=1$, etc., producing the following binary sequence:
@@
Other examples with $\mu=2$ can be constructed using logical operators. For instance, applying the AND-function with the rules
@@
to the initial pair 0 10\,1 produces the sequence
$s_{2}=0\,1\,0\,0\,0\,0\,0\,\ldots$
Similarly, applying the OR-function
1 11,1 $\rightarrow 1,\quad 1\,0\rightarrow 1,\quad 0\,1\rightarrow 1,\quad 0\,0\rightarrow 0$
to the same initial configuration 0 10\,1 results in
$s_{3}=0\,1\,1\,1\,1\,1\,1\,\ldots$
In contrast, applying the XOR-function
1 10,1 01,0 11,0 $\rightarrow 0,\quad 1\,0\rightarrow 1$$\quad$ $\rightarrow 1,\quad 0\,$$\rightarrow 0$
to the initial state $a_{1}=0,a_{2}=1$ yields
@@
with period $T_{4}=3$. However, using the same rule with the initial configuration $a_{1}=0,$ $a_{2}=0$ results in a constant sequence $a_{k}=0$ for all $k>2$. These examples illustrate that the resulting behavior depends not only on the rule but also on the initial conditions.
For any finite memory $\mu<\infty$, all sequences generated in this framework are eventually periodic. That is, for any rule and initial condition, the sequence will eventually settle into a repeating pattern (also referred to as a motif or loop). The maximum possible period for a sequence with memory $\mu$ is
@@
As $\mu$ increases, the sequences may resemble aperiodic ones, making them useful for computing pseudorandom numbers to be applied in Monte Carlo simulations and cryptographic systems (Etzion, ). Truly aperiodic behavior can only arise when $\mu\rightarrow\infty$, or when the alphabet is infinite (e.g., using real-valued states).
@@
where $a_{i}\in\{0,1\}$ for all $i=1,2,\ldots,\mu$ (see Figure 1). Starting from an initial configuration, the rule is applied recursively to generate the digits of the sequence as follows:
@@
Figure 1: Two generating rules with memory $\mu=3$. (A) Rule 150, whose binary representation is $10010110$. This corresponds to the following assignments for each 3-tuple (ordered from $111$ to $000$): $11\rightarrow 1, 11$$\rightarrow 0, 1010$$1\rightarrow 0$, $0\rightarrow 1, 0110$$1\rightarrow 0$, $\rightarrow 1, 001100$$\rightarrow 1$, $000\rightarrow 0$. The application of this rule to four different initial triplets is shown on the right of panel (A). As observed, the resulting sequences are periodic with periods 1, 4, 1, and 2 (from top to bottom). The dynamics of this rule can be visualized as a directed graph, shown at the bottom. Each node represents a possible 3-tuple; a directed edge from node $x$ to node $y$ exists if applying the rule to $x$ yields the last digit of $y$. This graph is not connected and does not contain a Hamiltonian path. (B) Rule 45, with binary representation $00101101$, corresponds to the truth table shown on the left of panel B. In contrast to Rule 150, applying Rule 45 to any of the $2^{3}=8$ initial triplets yields the same cyclic sequence of maximum period $T_{\text{max}}=8$, namely, the de Bruijn sequence $s_{6}=00010111$. This rule is thus a de Bruijn rule. The associated graph, built as described above, is a de Bruijn graph because it contains a Hamiltonian path through all $2^{3}$ nodes.
For example, when $\mu=3$, there are $2^{2^{3}}=256$ possible rules—corresponding to Wolfram’s elementary 1dCA rules. Figure 1A shows the XOR rule for $\mu=3$, represented by the binary rule string $10010110$, which corresponds to rule number 150 in Wolfram’s classification. Applying this rule to the initial configuration 0 1 00\,1\,0 yields
@@
In general, for a given $\mu=1,2,3,\ldots$, the total number of possible generating rules is $C(\mu)=2^{2^{\mu}}$ (this number corresponds with the Fermat number minus 1 A000215 ), and the maximum number of binary sequences generated from these rules is $2^{2^{\mu}+\mu}$. This number grows super-exponentially with $\mu$, making exhaustive analysis computationally infeasible for large $\mu$. For instance, when $\mu=10$, there are approximately $10^{308}$ rules, each of which must be applied to 1024 different initial configurations.
This raises the following fundamental question: given a memory value $\mu$ and an initial configuration, is it possible to predict the asymptotic pattern and period $T$ of the resulting sequence? While this is tractable for small $\mu$ via exhaustive enumeration, alternative methods are necessary for large $\mu$ due to the sheer size of the rule space.
It is worth noting that the 0-dimensional cellular automata (0dCA) with memory, as defined above, are equivalent to sequences generated by a Non-Linear Feedback Shift Register (NLFSR). This type of sequence has been extensively studied, and a vast body of literature exists on the subject Golomb ; Fredricksen ; Gao ; Zhao . In particular, significant attention has been devoted to the characterization and computation of shift register sequences with maximum period—commonly known as de Bruijn sequences deBruijn ; Ralston ; Huang ; Miros . As we will show in the next section, there exists a one-to-one correspondence between de Bruijn sequences and their generating rules, which we will refer to as de Bruijn rules.
@@ 3 Rules That Generate Maximum Period Sequences: De Bruijn Rules
For a given memory length $\mu$, the maximum period that a generated binary sequence can attain is $T_{\text{max}}=2^{\mu}$. Naturally, the minimum period is 1, corresponding to sequences that converge to fixed points (e.g., either 0 or 1). For small values of $\mu$, it is feasible to generate all possible sequences and compute their corresponding periods. For $\mu=4$, Table 1 shows the number of rules that generate sequences with periods in between 1 and $2^{4}$ for each of the $2^{4}$ initial sequence of length $4$. Note that the distributions of periods are symmetric to middle rows and that the maximum period occurs for only 16 rules, independently of the initial sequence. The proportion of rules as a function of the maximum period of the sequences that each of the $2^{2^{4}}$ generates over all the initial sequences is depicted in Figure 2 (also for the case $\mu=4$). It is worthy to remark the nonmonotonous shape of this distribution that exhibits three relative maxima at periods $T=1$, $T=3$, and $T=5$. For $T>5$, it seems to decrease exponentially.
@@
Sequences with a maximum period, known as de Bruijn sequences, are of particular interest. These are cyclic sequences of length $2^{\mu}$ in which every possible binary substring of length $\mu$ appears exactly once. For small values of $\mu$, it is straightforward to generate all possible de Bruijn sequences. There exists a unique de Bruijn sequence for $\mu=2$, namely, $0011$. For $\mu=3$, there are two such sequences: $00010111$ and $00011101$ that are generated by rules 45 and 75, respectively (rule 45 is graphically described in Figure 1B).
@@
Since these sequences are cyclic, their representations are not unique; by convention, we select the lexicographically least sequence (i.e., the smallest in binary order and decimal number) as the representative of the equivalence class. Specially significative is the lexicographically least sequence for each memory $\mu$ that is referred to as the granddaddy in Knuth .
Table 1: Number of rules that generates sequences with each of the possible periods (columns) that can appear, from 1 to 16, for $\mu=4$, applied to the $2^{4}$ initial sequences (rows). Notice that the total number of rules for this $\mu$-value is $2^{2^{4}}$ = 65,536, which equal the sum of each row. As it can be observed, the distributions of periods are symmetric to middle rows. Note that the maximum period occurs at the last column ($T=16$) for 16 rules and for every initial condition.
Initial Sequence 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
$00$ 36,096 3304 4944 4734 5836 3758 2472 1846 1120 758 332 144 80 48 48 16
0001 12,912 6608 9888 9468 9624 6492 3920 2924 1728 1068 440 208 96 80 64 16
0010 11,800 8976 11,780 8457 8976 5697 3552 2803 1634 995 412 200 94 80 64 16
0011 15,368 5320 9490 9252 9296 6397 3702 2981 1756 1067 444 207 96 80 64 16
0100 14,568 5524 11,588 8860 9190 5851 3570 2834 1672 1011 414 200 94 80 64 16
$01$ 10,520 16,384 8458 6494 8542 5755 3620 2572 1438 919 406 192 80 76 64 16
0110 12,032 6428 12,398 9128 9438 6043 3486 2975 1714 1019 422 199 94 80 64 16
0111 21,376 4956 7416 8125 8242 5637 3580 2641 1648 1025 426 208 96 80 64 16
1000 21,376 4956 7416 8125 8242 5637 3580 2641 1648 1025 426 208 96 80 64 16
1001 12,032 6428 12,398 9128 9438 6043 3486 2975 1714 1019 422 199 94 80 64 16
$10$ 10,520 16,384 8458 6494 8542 5755 3620 2572 1438 919 406 192 80 76 64 16
1011 14,568 5524 11,588 8860 9190 5851 3570 2834 1672 1011 414 200 94 80 64 16
1100 15,368 5320 9490 9252 9296 6397 3702 2981 1756 1067 444 207 96 80 64 16
1101 11,800 8976 11,780 8457 8976 5697 3552 2803 1634 995 412 200 94 80 64 16
1110 12,912 6608 9888 9468 9624 6492 3920 2924 1728 1068 440 208 96 80 64 16
$11$ 36,096 3304 4944 4734 5836 3758 2472 1846 1120 758 332 144 80 48 48 16
Figure 2: Bar chart of the proportion of rules as a function of the maximum period of the sequences that each of them generates over the initial sequences for a memory $\mu=4$. Note that $T=16$ has the lower occurrence; as a matter of fact, there are $2^{2^{3}-4}$ de Bruijn rules, which result in a relative frequency of $16/65,536\approx 0.0002$.
@@
The number of binary de Bruijn sequences for a given $\mu$ is given by $C(\mu)=2^{2^{\mu-1}-\mu}$ (deBruijn, ; A016031, ). This number grows also superexponentially with $\mu$ and reaches extremely large values even for moderate $\mu$. For example, for $\mu=10$, $C(10)\approx 10^{151}$. Generating all de Bruijn sequences is a challenging combinatorial problem that remains unsolved in full generality, though various efficient construction algorithms have been proposed (see, for instance, Hall ; deBruijnproject ).
@@
For $\mu=4$, there are $C(4)=2^{2^{3}-4}=16$ de Bruijn rules and their associated de Bruijn sequences. As it can be seen, the granddaddy sequence is $0000100110101111$. The corresponding granddaddy rule is $3825$, with binary representation $0000111011110001$. For $\mu=5$, the number of de Bruijn sequences is $C(5)=2^{2^{4}-5}=2048$. The granddaddy rule in this case is $00001100111111101111001100000001$ (whose decimal representation is 218,034,945), which generates the granddaddy sequence $00000100011001011011101010011111$.
Table 2: Correspondence betweeen de Bruijn rules and de Bruijn sequences for $\mu=1,2,3$ and $4$. The table also shows the Evil Odd Number that divides each de Bruijn rule, leaving the remainder $\phi(\mu)$. The Evil Odd Numbers that correspond to de Bruijn rules for $\mu=5$ are presented in Appendix B.
@@
4 43 255 10,965 0010101011010101 0000101001101111
4 57 255 14,535 0011100011000111 0000110111100101
4 65 255 16,575 0100000010111111 0000111101001011
4 71 255 18,105 0100011010111001 0000100111101011
4 75 255 19,125 0100101010110101 0000101111010011
4 83 255 21,165 0101001010101101 0000101100111101
4 85 255 21,675 0101010010101011 0000111101011001
4 89 255 22,695 0101100010100111 0000110010111101
4 99 255 25,245 0110001010011101 0000101001111011
4 113 255 28,815 0111000010001111 0000111101100101